Honesty15:

The volume of a rectangular prism is represented by the function x3 + 9x2 + 6x − 16. The length of the box is x + 2, while the height is x + 8. Find the expression representing the width of the box.

jhonyy9:

you need to know formula for volume of a rectangular prisme - do you know ?

supie:

\[V=volume\]\[L=length\]\[W=width\]\[H=hight\] \[V=L*W*H\] Values of Length and width \[L=(x+2)\]\[H=(x+8)\] Roots\[x = -8\]\[x = -2\]\[x = 1\] \[V=x^3 + 9x^2 + 6x − 16\] Using what is above ^ <that, \[V= x^3 + 9x^2 + 6x − 16 = (x+8)(x+2)*(x-1)\] I think that's right 🤔

darkknight:

L = (x+2) H=(x+8) W = ??? V = x^3+9x^2+6x-16 L(H)W = (x+8)(x+2)W v = LHW x^3+9x^2+6x-16 = (x+8)(x+2)(W) W = (x^3+9x^2+6x-16)/((x+8)(x+2)) W = ???